Direct Labor Variances

The following data relate to labor cost for production of 5,900 cellular telephones:

Actual: 3,990 hrs. at $16.10 $64,239
Standard: 3,930 hrs. at $16.40 $64,452

a. Determine the direct labor rate variance, direct labor time variance, and total direct labor cost variance. Enter all values as a positive number and then identify the variance (e.g. Favorable or Unfavorable) in the adjacent column.

Rate variance $
Time variance $
Total direct labor cost variance $

b. The employees may have been less-experienced or poorly trained, thereby resulting in a labor rate than planned. The lower level of experience or training may have resulted in efficient performance. Thus, the actual time required was than standard.
Answer #1
a
Rate variance -1197 Favorable =3990*(16.1-16.4)
Time variance 984 Unfavorable =16.4*(3990-3930)
Total direct labor cost variance -213 Favorable =64239-64452
b
The employees may have been less experienced or poorly trained, thereby resulting in a lower labor rate than planned. The lower level of experience or training may have resulted in less efficient performance. Thus, the actual time required was more than standard.Fortunately, the lost efficiency is offset by the lower labor rate.